OpenStudy (anonymous):

what is the solution of the proportion 5/x + 4 = 3/ x - 2

OpenStudy (anonymous):

5/(x+4) =3/(x-2) 5(x-2) = 3(x+4)

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

Do u know this? If a/b = c/d then ad =cb

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

Then u got this also right? 5/(x+4) =3/(x-2) 5(x-2) = 3(x+4)

OpenStudy (anonymous):

is the answer 10

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

wat bout 9 cuz i just did it and it give me 9

OpenStudy (anonymous):

5/(x+4) =3/(x-2) 5(x-2) = 3(x+4) 5x-10=3x+12 2x=22 x=11

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

got your mistake?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

welcome